# Tokenization Falling Short: hTe Cusre of Tkoeniaztion

### v1

@@ Abstract
Language models typically tokenize raw text into sequences of subword identifiers from a predefined vocabulary, a process inherently sensitive to typographical errors, length variations, and largely oblivious to the internal structure of tokens—issues we term the curse of tokenization. In this study, we delve into these drawbacks and demonstrate that large language models (LLMs) remain susceptible to these problems. This study systematically investigates these challenges and their impact on LLMs through three critical research questions: (1) complex problem solving, (2) token structure probing, and (3) resilience to typographical variation. Our findings reveal that scaling model parameters can mitigate the issue of tokenization; however, LLMs still suffer from biases induced by typos and other text format variations. Our experiments show that subword regularization such as BPE-dropout can mitigate this issue. We will release our code and data to facilitate further research.
@@
Yekun Chai∗♠  Yewei Fangthanks: Equal contribution and shared co-first authorship.  Qiwei Peng Xuhong Li Baidu   ModelBest   UCPH {chaiyekun,ywfcast1e}@gmail.com
@@ 1 Introduction
RQ3: Typographical Variation5). We designed a robust set of evaluation benchmarks termed TKEval on top of various datasets such as MMLU, TruthfulQA, GSM8K, and HumanEval, covering diverse tasks and linguistic phenomena. These benchmarks allow us to systematically test and analyze the LLMs’ resilience to tokenization.
Our findings highlight that while scaling model parameters can enhance the robustness of tokenization, LLMs still suffer from biases introduced by typographical errors and text format variations. We demonstrate the persistent nature of these tokenization challenges.
Contribution    1) We provide a comprehensive analysis of the problem known as the curse of tokenization, detailing its impact on language model performance and introducing systematic evaluation benchmarks to assess these issues.
2) By evaluating various scales of LLMs, including LLama3, Mistral, and GPT-4 families, across thirteen distinct tasks, we demonstrate that even state-of-the-art models struggle with handling typographical variations. Specifically, LLMs exhibit greater sensitivity to character-level variations compared to subword-level variations.
3) We demonstrate that regularized tokenization approaches, such as BPE-dropout with moderate drop rates, can enhance the model’s resilience to the discussed issues.
4) To facilitate further research, we will release our evaluation code and benchmarks, enabling the research community to build upon our findings and develop more robust models.
@@ 2.1 Tokenization
Tokenization Approach Conventional language models Radford et al. (2018); Chai et al. (2020) typically tokenize input text into a sequence of tokens by splitting it into smaller subwords. Traditional tokenization approaches include frequency-based methods such as Byte Pair Encoding (BPE; Sennrich et al., 2016) and probability-based methods like WordPiece Schuster and Nakajima (2012). BPE merges tokens based on bigram frequency, relying on subword pair co-occurrence to greedily merge neighboring pairs. In contrast, WordPiece can be viewed as a language-modeling based BPE variant. It select the unit pair that maximizes the bigram likelihood of training data at utmost, rather than choose the most frequent pair.
@@
Tokenization-Free Approach Tokenization approaches often suffer from the vocabulary bottleneck, where there is a trade-off between vocabulary size and diverse language coverage in multilingual scenarios. To address this issue, Rust et al. (2023); Chai et al. (2024) introduced a tokenization-free approach that renders raw text as visual text images for masked language modeling and autoregressive pre-training. This method demonstrates robust multilingual generalization capabilities compared to subword tokenization approaches.
@@ 2.2 Perturbation Probing
Several studies have investigated the behavior of language models under input perturbations at various levels, including character-level Nishino et al. (2019), subword-level Abdou et al. (2022), and word-level Sinha et al. (2021) scrambling. Despite these efforts, intrinsic evaluations of perturbing LLM inputs remain under-explored.
@@ 3.1 Anagram Task
Task Description and Settings   The anagram task tests the model’s ability to unscramble a sequence of jumbled characters to form a valid word. This task evaluates the model’s handling of surface-form manipulations and its understanding of char-level compositions within a word. The complexity arises from the need to identify potential word candidates from mixed characters and reassemble them correctly. We present a task example in §A.1. Specifically, we include two tasks:
@@
  • Word Unscrambling (WU) Srivastava et al. (2022) – The model is given a randomly scrambled word, and must recover the original word (e.g., “nad” $\rightarrow$ “and”).
@@
Results and Analysis   The experimental results reveal that larger models demonstrate better performance on the anagram task, yet they remain susceptible to tokenization errors. Specifically, models struggled with longer anagrams (see Figure 3) or those containing uncommon letter combinations, indicating that while scaling improves token recognition, inherent tokenization flaws persist.
Figure 3: Violin plot illustrating the relationship between the length of scrambled words and the Exact Match (EM) score of Llama3-8B and Llama3-70B on the word unscrambling task under one-shot evaluation. The models tend to correctly reorder anagrams of shorter lengths, while struggling with longer words.
The violin plot in Figure 3 highlights the performance differences between Llama3-8B and Llama3-70B on the word unscrambling task across various word lengths. Notably, Llama3-70B consistently outperforms Llama3-8B, especially in the 0-6 and 6-12 character buckets, where it shows higher and denser EM scores. This trend indicates that as the model parameter size increases from 8B to 70B, the ability to accurately reorder scrambled words improves. However, both models struggle with longer words (12-18 characters), though Llama3-70B maintains a slight edge.
Our results indicate several key trends. Firstly, the performance improves significantly as model size scales from 8B to 70B parameters, as evidenced by the performance of Llama3-70B compared to Llama3-8B AI@Meta (2024). Secondly, we observe that while the dense Mistral-7B model performs poorly, the sparse Mixtral-8x7B model (an MoE sparse model) shows improved performance due to its parameter size scaling. Lastly, GPT-4 turbo, a much more powerful model, achieves state-of-the-art results, clearly outperforming all other models across all shot conditions. This sensitivity underscores the need for more robust tokenization that can handle typographical variations without degrading performance.
@@ 3.2 Mathematical Language () Comprehension
Table 1: Few-shot results on Identifying Math Theorems, with exact match (EM) scores reported as percentages. a[Srivastava et al., 2022]
Results and Analysis    Our evaluation results of various models are shown in Table 1. The results demonstrate that while larger models generally perform better on LaTeX-formatted mathematical content, the relationship between the number of in-context examples and model performance is not linear. The Llama3-70B model consistently outperformed other models, achieving a score of 62.26% in the zero-shot setting and improving to 79.25% with one-shot learning. However, additional in-context examples led to fluctuating performance, with scores of 69.81% in the two-shot setting and 71.70% in the three-shot setting. This indicates that increasing the number of in-context demonstrations does not consistently enhance performance and may lead to variability depending on the specific examples provided.
@@ 4.1 Intra-Token Probing
Task Description and Settings   To measure the capability of LLMs, we devise intra-token probing tasks related to length, case, and counting problems. These tasks evaluate the model’s performance on the internal structure of tokens or word, specifically including four tasks:
@@
Results and Analysis   We evaluated CC, NC, NCR, and CCV tasks across different models and shot settings. The results are presented in Figure 4.
@@ 4.2 Inter-Token Probing
Task Description and Settings   To evaluate the capabilities of LLMs in understanding and manipulating relationships between multiple tokens, we devised inter-token probing tasks. These tasks focus on identifying common patterns and sequences across tokens, and they assess the model’s ability to recognize and process such relationships. Specifically, we include three tasks:
@@
Results and Analysis   The results are presented in Figure LABEL:fig:inter-token-probe. The results for CS tasks indicate that larger models, such as Llama3-70B and GPT-4 Turbo, perform significantly better than smaller models. GPT-4 Turbo achieves the highest accuracy across all shot settings, demonstrating the model’s robustness in identifying continuous substrings. Notably, Llama3-70B also shows strong performance, particularly in the three-shot setting. Sparse models like Mixtral-8x7B exhibit notable improvements compared to dense models, highlighting the effectiveness of sparse architectures in handling complex token relationships.
@@ 5 RQ3: Typographical Variation
Task Description and Settings   The primary goal of these tasks is to assess the LLMs’ resilience to typographical errors at both the character and token levels, examining whether these models can maintain semantic understanding when faced with such perturbations. For character-level typographical variation, we employed $n$-gram shuffling within word boundaries (with $n$ set to 2, 3, 5) with a 50% probability, and $n$-gram noise, which involve adding, deleting, and replacing characters, spaces, and punctuation marks to simulate spelling noise. This corruption occurs with a 30% probability, including insertion, deletion, or substitution operations. Token-level typographical variation was introduced by shuffling tokens within $n$-grams of sizes 2, 3, and 5, with a 50% probability of permutation or typo generation, similar to the character-level method. We include four tasks:
@@
Results and Analysis   Figures LABEL:fig:exp3 presents a comprehensive analysis of the impact of typographical variations on various LLMs, specifically focusing on character-level and token-level perturbations across different $n$-gram sizes ($n$=2, 3, 5). The evaluation covers a range of datasets, including TruthfulQA, MMLU, GSM8K (5-shot), and HumanEval.
@@ 6 Does BPE-dropout Matter?
To further enhance the robustness of our models, we explore regularized tokenization approach, BPE-dropout Provilkov et al. (2020), which randomly drops BPE merges during tokenization. This technique allows text sequences to be tokenized in more diverse ways, promoting robustness to various token combinations and increasing the likelihood of encountering smaller tokens. Intuitively, this diversity benefits the model’s understanding of internal token structures.
Training Setup   For the training data, we synthesized a dataset consisting of 111k examples specifically designed for RQ2. More detailed training settings are provided in Appendix B.3.
Results and Analysis   Figure 6 shows the impact of varying BPE-dropout rates on the Mistral-7B model’s performance across multiple $K$-shot settings and tasks. The baseline performance, with a dropout rate of $p=0$, shows robust results across several tasks, particularly in CS and CC. These tasks are relatively straightforward, involving simple token manipulations that do not significantly challenge the model’s capacity to generalize from zero-shot to few-shot scenarios. The high performance in these tasks suggests that the Mistral-7B model, even without BPE-dropout, is adept at handling simpler token relationships. However, it is important to note that the baseline performance does not uniformly extend to more complex tasks.
@@
Our analysis reveals that higher dropout rates ($p=0.6$ and $p=0.8$) exhibit relatively lower performance across most tasks. This decline can be attributed to insufficient training, as the dataset was trained for only five epochs. The higher dropout rates introduce greater tokenization variation, which necessitates additional training compute to achieve convergence. The lack of adequate training epochs likely hindered these models from fully leveraging the potential benefits of higher BPE-dropout rates. We provided detailed training analysis in Appendix §D.
@@ B.3 Post-Training Details
Training Details
We employ the AdamW optimizer Loshchilov and Hutter (2017) with the hyperpameters of $\beta_{1}=0.9$, $\beta_{2}=0.95$. The peak learning rate is set to 5e-5, and the minimum learning rate is set to 1e-6. The learning rate warms up during the first 10% of training steps and then decays with a cosine scheduler.
Given the difference in data distribution resulting from BPE-dropout, we post-train Mistral-7B Jiang et al. (2023) on the training split for 5 epochs with a global batch size of 16. Following the pre-training recipe, we concatenate all sequences and then chunk them into fixed context lengths of 4096 for our autoregressive post-training. We conduct data shuffling within the same epochs.
@@ C.2 Typographical Variation Task (RQ3)
We performed character-level corruption within word boundaries, ensuring that the positions of punctuation marks or spaces remained unchanged after corruption. Even when there were not enough characters to form a complete $n$-gram, corruption was still applied.
@@
We converted the input prompts into tokens using tokenizers from different model families. Corruption was then applied within an $n$-gram range, without regard to word boundaries. This approach simulates token-level typographical variations, challenging the models to handle disruptions at the token level.
Figure 7: Example of a char-level permutation(5-gram) prompt from HumanEval.
Permutation
For permutation task, we randomly shuffle tokens or characters within an $n$-gram range with a 50% probability, as illustrated in Figure 7. We evaluate by using $n$-gram range from 2 to 5, in which various $n$-gram levels could assess the model’s performance under varying degrees of corruption.
Noise Injection
Figure 8: Example of token-level noise (tri-gram) injected prompt from GSM8K. Red, Green, and Gray denote replacement, insertion, and deletion respectively.
When performing replacing, we replace a token or character in the current $n$-gram with a token or character randomly selected from the whole vocabulary. Figure 7 illustrates our approach of injecting noise to tokens using a tri-gram granularity.

### v2

@@ Abstract
Language models typically tokenize raw text into sequences of subword identifiers from a predefined vocabulary, a process inherently sensitive to typographical errors, length variations, and largely oblivious to the internal structure of tokens—issues we term the curse of tokenization. In this study, we delve into these drawbacks and demonstrate that large language models (LLMs) remain susceptible to these problems. This study systematically investigates these challenges and their impact on LLMs through three critical research questions: (1) complex problem solving, (2) token structure probing, and (3) resilience to typographical variation. Our findings reveal that scaling model parameters can mitigate the issue of tokenization; however, LLMs still suffer from biases induced by typos and other text format variations. Our experiments show that subword regularization such as BPE-dropout can mitigate this issue. We release our evaluation code and data at https://github.com/FloatAI/TKEval.
@@
Yekun Chai∗♠  Yewei Fangthanks: Equal contribution and shared co-first authorship.  Qiwei Peng Xuhong Li Baidu   ModelBest   University of Copenhagen {chaiyekun,ywfcast1e}@gmail.com
@@ 1 Introduction
RQ3: Typographical Variation5). We designed a robust set of evaluation benchmarks on top of various datasets such as MMLU, TruthfulQA, GSM8K, and HumanEval, covering diverse tasks and linguistic phenomena. These benchmarks allow us to systematically test and analyze the LLMs’ resilience to tokenization.
Our findings highlight that while scaling model parameters can enhance the robustness to tokenization, LLMs still suffer from biases introduced by typographical errors and text format variations. We demonstrate the persistent nature of these tokenization challenges.
Contribution
To conclude, our main contributions are as follows:
  1. 1.
    We provide a comprehensive analysis of the problem known as the curse of tokenization, detailing its impact on language model performance and introducing systematic evaluation benchmarks to assess these issues.
  2. 2.
    By evaluating various scales of LLMs, including LLama3, Mistral, and GPT-4 families, across thirteen distinct tasks, we demonstrate that even state-of-the-art models struggle with handling typographical variations. Specifically, LLMs exhibit greater sensitivity to character-level variations compared to subword-level variations.
  3. 3.
    We demonstrate that regularized tokenization approaches, such as BPE-dropout with moderate drop rates, can enhance the model’s resilience to the discussed issues.
@@ 2.1 Tokenization
Tokenization Approach
Conventional language models Radford et al. (2018); Jozefowicz et al. (2016); Brown et al. (2020) typically tokenize input text into a sequence of tokens by splitting it into smaller subwords. Traditional tokenization approaches include frequency-based methods such as Byte Pair Encoding (BPE; Sennrich et al., 2016) and probability-based methods like WordPiece Schuster and Nakajima (2012). BPE merges tokens based on bigram frequency, relying on subword pair co-occurrence to greedily merge neighboring pairs. In contrast, WordPiece can be viewed as a language-modeling based BPE variant. It select the unit pair that maximizes the bigram likelihood of training data at utmost, rather than choose the most frequent pair.
@@
Tokenization-Free Approach
Tokenization approaches often suffer from the vocabulary bottleneck, where there is a trade-off between vocabulary size and diverse language coverage in multilingual scenarios. To address this issue, Rust et al. (2023) and Chai et al. (2024) introduced a tokenization-free approach that renders raw text as visual text images for masked language modeling and autoregressive pre-training. This method demonstrates robust multilingual generalization capabilities compared to subword tokenization approaches.
@@ 2.2 Perturbation Probing
Several studies have investigated the behavior of language models under input perturbations at various levels, including character-level Nishino et al. (2019), subword-level Abdou et al. (2022), and word-level Sinha et al. (2021); Peng et al. (2023) scrambling. Despite these efforts, intrinsic evaluations of perturbing LLM inputs remain under-explored.
@@ 3.1 Anagram Task
Task Description and Settings
The anagram task tests the model’s ability to unscramble a sequence of jumbled characters to form a valid word. This task evaluates the model’s handling of surface-form manipulations and its understanding of char-level compositions within a word. The complexity arises from the need to identify potential word candidates from mixed characters and reassemble them correctly. We present a task example in §A.1. Specifically, we include two tasks:
@@
  • Word Unscrambling (WU; Srivastava et al., 2022) – The model is given a randomly scrambled word, and must recover the original word (e.g., “nad” $\rightarrow$ “and”).
@@
Results and Analysis
The experimental results reveal that larger models demonstrate better performance on the anagram task, yet they remain susceptible to tokenization errors. Specifically, models struggled with longer anagrams (see Figure 3) or those containing uncommon letter combinations, indicating that while scaling improves token recognition, inherent tokenization flaws persist.
Figure 3: The relationship between the length of scrambled words and the Exact Match (EM) score of Llama3-8B and Llama3-70B on the word unscrambling task under one-shot evaluation. The models tend to correctly reorder anagrams of shorter lengths, while struggling with longer words.
Figure 3 highlights the performance differences between Llama3-8B and Llama3-70B on the word unscrambling task across various word lengths. Notably, Llama3-70B consistently outperforms Llama3-8B, especially in the 0-6 and 6-12 character buckets. This trend indicates that as the model parameter size increases from 8B to 70B, the ability to accurately reorder scrambled words improves. However, both models struggle with longer words (12-18 characters), though Llama3-70B maintains a moderate edge.
Our results shown in Figure 2 indicate several key trends. Firstly, the performance improves significantly as model size scales from 8B to 70B parameters, as evidenced by the performance of Llama3-70B compared to Llama3-8B AI@Meta (2024). Secondly, we observe that while the dense Mistral-7B model performs poorly, the sparse Mixtral-8x7B model (an MoE sparse model) shows improved performance due to its parameter size scaling. Lastly, GPT-4 turbo, a much more powerful model, achieves state-of-the-art results, clearly outperforming all other models across all shot conditions. This sensitivity underscores the need for more robust tokenization that can handle typographical variations without degrading performance.
@@ 3.2 Mathematical Language () Comprehension
Table 1: Few-shot results on Identifying Math Theorems, with exact match scores reported as percentages. a refers to results taken from Srivastava et al. (2022)
Results and Analysis
Our evaluation results of various models are shown in Table 1. The results demonstrate that while larger models generally perform better on LaTeX-formatted mathematical content, the relationship between the number of in-context examples and model performance is not linear. The Llama3-70B model consistently outperformed other models, achieving a score of 62.26% in the zero-shot setting and improving to 79.25% with one-shot learning. However, additional in-context examples led to fluctuating performance, with scores of 69.81% in the two-shot setting and 71.70% in the three-shot setting. This indicates that increasing the number of in-context demonstrations does not consistently enhance performance and may lead to variability depending on the specific examples provided.
@@ 4.1 Intra-Token Probing
Task Description and Settings
To measure the capability of LLMs, we devise intra-token probing tasks related to length, case, and counting problems. These tasks evaluate the model’s performance on the internal structure of tokens or word, specifically including four tasks:
@@
Results and Analysis
We evaluated CC, NC, NCR, and CCV tasks across different models and shot settings. The results are presented in Figure 4.
@@ 4.2 Inter-Token Probing
Task Description and Settings
To evaluate the capabilities of LLMs in understanding and manipulating relationships between multiple tokens, we devised inter-token probing tasks. These tasks focus on identifying common patterns and sequences across tokens, and they assess the model’s ability to recognize and process such relationships. Specifically, we include three tasks:
@@
Results and Analysis
The results are presented in Figure LABEL:fig:inter-token-probe. The results for CS tasks indicate that larger models, such as Llama3-70B and GPT-4 Turbo, perform significantly better than smaller models. GPT-4 Turbo achieves the highest accuracy across all shot settings, demonstrating the model’s robustness in identifying continuous substrings. Notably, Llama3-70B also shows strong performance, particularly in the three-shot setting. Sparse models like Mixtral-8x7B exhibit notable improvements compared to dense models, highlighting the effectiveness of sparse architectures in handling complex token relationships.
@@ 5 RQ3: Typographical Variation
Task Description and Settings
The primary goal of these tasks is to assess the LLMs’ resilience to typographical errors at both the character and token levels, examining whether these models can maintain semantic understanding when faced with such perturbations. For character-level typographical variation, we employed $n$-gram shuffling within word boundaries (with $n$ set to 2, 3, 5) with a 50% probability, and $n$-gram noise, which involve adding, deleting, and replacing characters, spaces, and punctuation marks to simulate spelling noise. This corruption occurs with a 30% probability, including insertion, deletion, or substitution operations. Token-level typographical variation was introduced by shuffling tokens within $n$-grams of sizes 2, 3, and 5, with a 50% probability of permutation or typo generation, similar to the character-level method. We include four tasks:
@@
Results and Analysis
Figures LABEL:fig:exp3 presents a comprehensive analysis of the impact of typographical variations on various LLMs, specifically focusing on character-level and token-level perturbations across different $n$-gram sizes ($n$=2, 3, 5). The evaluation covers a range of datasets, including TruthfulQA, MMLU, GSM8K (5-shot), and HumanEval.
@@ 6 Does BPE-dropout Matter?
To further enhance the robustness of LLMs, we explore regularized tokenization approach, BPE-dropout Provilkov et al. (2020), which randomly drops BPE merges during tokenization. This technique allows text sequences to be tokenized in more diverse ways, promoting robustness to various token combinations and increasing the likelihood of encountering smaller tokens. Intuitively, this diversity benefits the model’s understanding of internal token structures.
Training Setup
For the training data, we synthesized a dataset consisting of 111k examples specifically designed for RQ2. We employ the AdamW optimizer Loshchilov and Hutter (2017) with the hyperpameters of $\beta_{1}=0.9$, $\beta_{2}=0.95$. The peak learning rate is set to 5e-5, and the minimum learning rate is set to 1e-6. The learning rate warms up during the first 10% of training steps and then decays with a cosine scheduler.
Given the difference in data distribution resulting from BPE-dropout, we post-train Mistral-7B Jiang et al. (2023) on the training split for 5 epochs with a global batch size of 16. Following the pre-training recipe, we concatenate all sequences and then chunk them into fixed context lengths of 4096 for our autoregressive post-training. We conduct data shuffling within the same epochs. More detailed training settings are provided in Appendix B.3.
Results and Analysis
Figure 6 shows the impact of varying BPE-dropout rates on the Mistral-7B model’s performance across multiple $K$-shot settings and tasks. The baseline performance, with a dropout rate of $p=0$, shows robust results across several tasks, particularly in CS and CC. These tasks are relatively straightforward, involving simple token manipulations that do not significantly challenge the model’s capacity to generalize from zero-shot to few-shot scenarios. The high performance in these tasks suggests that the Mistral-7B model, even without BPE-dropout, is adept at handling simpler token relationships. However, it is important to note that the baseline performance does not uniformly extend to more complex tasks.
@@
Our analysis reveals that higher dropout rates ($p=0.6$ and $p=0.8$) exhibit relatively lower performance across most tasks. This decline can be attributed to insufficient training, as the dataset was trained for only five epochs. The higher dropout rates introduce greater tokenization variation, which necessitates additional training compute to achieve convergence. The lack of adequate training epochs likely hindered these models from fully leveraging the potential benefits of higher BPE-dropout rates. Moreover, we report the test-set performance across seven tasks over the course of BPE-dropout fine-tuning in Figure LABEL:fig:ft_epoch. We provided detailed training analysis in Appendix §D.
@@ 7 Conclusion
Acknowledgements
We would like to thank all anonymous reviewers for their insightful comments and feedback. Qiwei Peng is supported by DisAI - Improving scientific excellence and creativity in combating disinformation with artificial intelligence and language technologies, a project funded by European Union under the Horizon Europe, GA No. 101079164.
@@ C.2 Typographical Variation Task (RQ3)
Figure 10: Example of a char-level(5-gram) corrupted prompt from HumanEval. Red, Green, and Gray denote replacement, insertion, and deletion respectively.
Figure 11: Example of a token-level(tri-gram) corrupted prompt from GSM8K. Red, Green, and Gray denote replacement, insertion, and deletion respectively.
Permutation
For permutation task, we randomly shuffle tokens or characters within an $n$-gram range with a 50% probability. We evaluate by using $n$-gram range from 2 to 5, in which various $n$-gram levels could assess the model’s performance under varying degrees of corruption.
Noise Injection
When performing replacing, we replace a token or character in the current $n$-gram with a token or character randomly selected from the whole vocabulary.
@@
We performed character-level permutation within word boundaries, ensuring that the positions of punctuation marks or spaces remained unchanged after permutation. Even when there were not enough characters to form a complete $n$-gram, permutation was still applied. We ignore word boundaries when injecting character-level noise. Figure 10 illustrates our character-level corruption approach using a 5-gram granularity.
@@
We converted the input prompts into tokens using tokenizers from different model families. Corruption was then applied within an $n$-gram range, without regard to word boundaries. This approach simulates token-level typographical variations, challenging the models to handle disruptions at the token level. Figure 11 gives an example of how we apply such variations to an encoded prompt.